# Dielectric response in proteins: The proteotronics approach

### v1

@@ Abstract
The dielectric properties of proteins, particularly in their hydrated state, have been extensively studied. Numerous theoretical and experimental investigations have reported values of both the permittivity and the intrinsic dipole moments of specific proteins under well-defined hydration conditions. Since even approximate estimates of these properties are relevant from both fundamental and applied perspectives, we propose a user-friendly method to calculate the relative permittivity that can be readily integrated into proteotronics workflows. To validate the proposed approach, we compare the results with those obtained using a classical macroscopic method. The outcomes are consistent and contribute further insight into this long-debated issue.
@@
These computational approaches, including molecular dynamics (MD) and Monte Carlo (MC) simulations, are known to be time-consuming, and the results obtained do not always agree with the limited available experimental data. To address this complexity, coarse-grained methods have been developed more recently, based on a continuous distribution of dielectric permittivity within the protein Li et al. (2013); Grant et al. (2001). The central idea of this approach is to provide an effective description of the protein–solvent system: in surface regions and internal cavities, the effective permittivity is taken as an average between water and the protein residues, while within the protein core it approaches the permittivity of the dry residues. A key challenge is determining how to generate this spatial distribution. In the literature, this is often achieved by representing the protein as an equivalent sphere and continuously distributing permittivity values among its amino acids or atoms, based on measures such as the radius of gyration Li et al. (2013). However, this approach can yield inconclusive results for highly elongated proteins Li et al. (2013).
@@
Alternatively, we propose evaluating the effective permittivity by considering the number of nearest neighbors for each amino acid. This quantity is a key descriptor of the topological features of a protein’s specific conformation. Specifically, it is used to produce the protein representations through complex networks in a procedure is called proteotronicsAlfinito et al. (2015). Furthermore, it is commonly used to generate two-dimensional representations of the tertiary structure, such as contact maps. Based on this approach, permittivity values are assigned so that highly connected regions correspond to lower permittivity, while less connected regions are assigned higher values. This strategy provides a more accurate estimate of buried versus exposed regions, thereby addressing limitations reported in previous studies Li et al. (2013).
@@
To define nearest neighbors, we model the protein as a complex network in which nodes correspond to the centroids of amino acids Alfinito et al. (2015, 2011, 2010). For each node, nearest neighbors are defined as those nodes located within a distance less than $R_{C}$, the interaction radius. The choice of $R_{C}$ determines the number of nearest neighbors: very small values (< 4 Å) produce a disconnected network, where most nodes have only 0–1 neighbors, whereas very large values result in a fully connected network, with each amino acid linked to all others Alfinito et al. (2008, 2015, 2011). Neither of these extremes is of interest. In this study, we set $R_{C}=6$Å which captures the geometric regularity of proteins Simonson (2003); Alfinito et al. (2008, 2010) and yields an average of approximately five nearest neighbors per node. Tests performed for $R_{C}$ values in the range 5–7 Å show no qualitative differences in the observed trends.
@@
Finally, an appropriate distribution function must be selected. In principle, different choices may lead to significantly different outcomes. To address this issue, we evaluated several candidate functions, seeking those that yield higher permittivity values for elongated proteins while ensuring consistency across structurally similar proteins. These functions are formulated using nearest-neighbor descriptors, including the maximum value ($L_{\rm max}$) and the mean value ($\langle L\rangle$) computed over the entire protein.
@@
The test functions are employed to assign a local permittivity value, $\epsilon(n)$, to each $n$-th amino acid. [^1] Hereafter, the term permittivity refers to the relative permittivity, i.e., the permittivity normalized to the vacuum permittivity. This quantity depends on both the intrinsic (dry) permittivity, $\epsilon_{i}(n)$ , and the amino acid’s position within the native protein structure. More specifically, $\epsilon(n)$ varies between a maximum value corresponding to that of the solvent under standard temperature and pressure (STP) conditions and $\epsilon_{i}(n)$, determined as described in Appendix B. We use three quite standard test functions, namely, linear, sigmoidal and exponential, whose
sharpness can be tuned via a control parameter (Appendix A). Increasing the sharpness enhances the contrast between regions characterized by high and low permittivity, facilitating the identification of hydrated (“wet”) regions associated with larger $\epsilon(n)$ values. The distinction between an internal low-permittivity region and an external high-permittivity region underlies many theoretical and computational studies Simonson and Brooks (1996); Simonson (2003, 2008); Warshel et al. (2006), as the effective dielectric response critically depends on this partitioning.
@@
For this reason, in this Letter we analyze a dataset consisting of 31 proteins, including both single-chain and complex structures, selected to span a range of shapes and sizes and to perform different functions. The aim is to assess whether the proposed approach satisfies the minimal consistency requirements discussed above. The protein selection is arbitrary and may be extended in future studies. Figure 1 presents the normalized permittivity values, $(k-\langle k\rangle)/\sigma$, where $\langle k\rangle$ denotes the mean and $\sigma$ the standard deviation computed over the dataset. The set of test functions considered captures the differences between elongated and approximately spherical proteins. In particular, elongated proteins systematically exhibit higher average permittivity values than spherical structures.
All test functions show a comparable trend, which suggests that it is not the chosen function itself that is important, but rather the permittivity distribution based on the coordination number. For illustration, Fig. 1 reports the dataset-averaged permittivity obtained using the linear distribution function, shown separately for elongated and spherical proteins.
@@
Within the framework of classical electrodynamics, the electrical polarizability of a collection of weakly interacting dipoles is determined by the ensemble-averaged dipole orientation induced by an external electric field, $E$. The derivation follows from the thermal angular distribution function describing dipole orientations
@@
Hydrated proteins cannot be regarded as freely rotating entities, as the surrounding solvent constrains their orientation in the presence of both external and internal electric fields. The native structure itself arises from hydration effects and results from the folding process, which typically positions hydrophilic residues toward the solvent-exposed exterior and hydrophobic residues within the protein core. Water molecules near the surface dynamically reorient in response to local electrostatic fields, thereby reducing the internal electric field generated by charged and polar residues and contributing to structural stability. Consequently, the dipole moment reported in Ref. Felder et al., 2007 reflects this solvent-mediated charge organization. In contrast, the dipole moment of the dry protein differs due to the absence of solvent-induced polarization and screening effects.
The dipole moment of a dry protein, denoted as $p_{0}$, can be estimated following the approach proposed by Song Song (2002). In that paper, the intrinsic polarizabilities of amino acids in proteins were evaluated using molecular dynamics simulations. The term intrinsic refers to a condition in which the dipole moment of each amino acid is assumed to be unaffected by external influences (e.g., solvent effects) and by interactions with other dipoles within the protein Song (2002). This approximation corresponds to the limiting case of a dry protein, where the dipole moments of individual amino acids are sufficiently small to be considered nearly independent, and solvent-induced polarization is absent. Within this framework, the relative dielectric constant of a protein can be evaluated as $k_{0}=\langle\epsilon(n)\rangle$, where, for each amino acid, $\epsilon(n)=\epsilon_{i}$, i.e., the intrinsic dielectric constant obtained from Ref. Song, 2002 (see Appendix B). The resulting value of $k_{0}$ is more or less the same across the selected proteins, say, $k_{0}\approx 1.5$ , in fine agreement with the outcomes of Ref. Simonson, 2003 (see also Appendix B for details).
A collection of dry proteins may be regarded as freely orientable in an external electric field, due to the absence of solvent-induced friction. Consequently, the dielectric response of the system, i.e., the dry susceptibility, $\chi_{0}$, can be described by Eq. 4. Recognizing that the dipole moment of a hydrated protein, $p$, depends on the corresponding dry dipole moment $p_{0}$, the expectation value of polarization under hydrated conditions can be evaluated following Eq. (2). In particular:
$\langle p^{\parallel}(T)\rangle=\int p\,\cos\theta\,\mathcal{P}_{0}(T)\,d\cos\theta.$ (5)
@@
Finally, we note that Eq. 8, yields an effective macroscopic susceptibility scaling as $\chi^{H}\propto p\,\Omega^{-n}\equiv\chi^{H}(n)$, with $n=1/2$. To assess whether the choice $n=1/2$ optimizes the agreement with the continuous microscopic models, Fig. 3 compares the previously introduced linear model with $\chi^{H}(n)$ evaluated for $n=0$, $0.5$, and $1$. Within the adopted approximations, the closest agreement is observed for $n=1/2$.
@@
In conclusion, the calculation and definition of permittivity in proteins remain subjects of ongoing debate and are known to depend sensitively on the adopted theoretical and computational framework (see, e.g., Ref. Warshel et al., 2006). In this paper, we propose a computationally efficient and user-friendly approach for estimating the effective macroscopic susceptibility of hydrated proteins. This formulation is motivated by the need to describe the orientational response of macromolecules with restricted rotational mobility arising from their interaction with a polar solvent. The proposed procedure is not intended to provide a complete or fully rigorous description. Rather, it offers a practical strategy that may complement more detailed molecular dynamics approaches, which, despite their accuracy, have not yet produced a comprehensive reference dataset of dielectric properties across diverse protein classes.
@@
In summary, within classical electrodynamics, a system of freely rotating dipoles yields the susceptibility given by Eq. 3. This expression is not directly applicable to hydrated proteins (dipole moment $p$), as it leads to susceptibility values significantly larger than those reported experimentally (see Table 1). The discrepancy arises from the fact that proteins in aqueous environments are not freely orientable: their rotational dynamics are constrained by solvent interactions, which in turn limit their dielectric response. By contrast, the susceptibility evaluated using Eq. 8 produces values qualitatively consistent with the expected range Simonson (2003)(see Table 1). Furthermore, the susceptibility associated with dry proteins (dipole moment $p_{0}$), denoted as $\chi_{0}$, falls within the range typically attributed to the internal regions of proteins Simonson (2003). We emphasize that the difference between $p$ and $p_{0}$ does not primarily originate from structural modifications of the three-dimensional conformation, but rather from changes in polarizability. The dipole moment $p$ characterizes a hydrated protein, where interfacial water molecules dynamically reorient and influence the distribution of surface charges, thereby altering the dielectric response, particularly for solvent-exposed residues. In contrast, $p_{0}$ corresponds to the limiting case of a dry protein, in which the native structure is ideally preserved while solvent-induced polarization and screening effects are absent. Within this interpretation, $p$ represents a hydration-dependent state derived from $p_{0}$, where the dominant variation concerns the effective permittivity. Accordingly, $p$ may be regarded as a function of $p_{0}$, and its expectation value can be estimated using the orientational probability distribution associated with the freely rotating dipole $p_{0}$. The results obtained using the two independent procedures show qualitative consistency, supporting the validity of the proposed framework. Nevertheless, further refinements are possible, particularly regarding the selection of test functions, the choice of the interaction radius, $R_{C}$ and the evaluation of dipole moments. Finally, we note that in physiological and experimental conditions proteins are generally not freely mobile. This observation suggests that additional investigation is warranted to clarify the definition and interpretation of electrical susceptibility in constrained biological environments.

### v2

@@ Abstract
The dielectric properties of proteins, particularly in their hydrated state, have been extensively studied. Numerous theoretical and experimental investigations have reported values of both the permittivity and the intrinsic dipole moments of specific proteins under well-defined hydration conditions. Since even approximate estimates of these properties are relevant from both fundamental and applied perspectives, we propose a ease-to-use method to calculate the relative permittivity that can be readily integrated into proteotronics workflows. To validate the proposed approach, we compare the results with those obtained using a classical macroscopic method. The outcomes are consistent and contribute further insight into this long-debated issue.
@@ I Introduction
These computational approaches, including molecular dynamics (MD) and Monte Carlo (MC) simulations, are known to be time-consuming Matyushov (2012); Amin and Küpper (2020); Hartl et al. (2022); Tenenbaum (2024); Agelii et al. (2025), and the results obtained do not always agree with the limited available experimental data . To address this complexity, coarse-grained methods have been developed more recently, based on a continuous distribution of dielectric permittivity within the protein Li et al. (2013); Grant et al. (2001). The central idea of this kind of approach (smooth permittivity Li et al. (2013)) is to provide an effective description of the protein–solvent system: in surface regions and internal cavities the effective permittivity is taken as an average between water and the protein residues, while within the protein core it approaches the permittivity of the dry residues. A key challenge is determining how to generate this spatial distribution. In the literature, this is often achieved by representing the protein as an equivalent sphere and continuously distributing permittivity values among its amino acids or atoms, based on measures such as the radius of gyration Li et al. (2013). However, this approach can yield inconclusive results for highly elongated proteins Li et al. (2013).
@@
Following the continuous method, we propose to evaluate the effective permittivity by evaluating the shape of the protein not by the radius of gyration, but rather by the number of nearest neighbors (coordination number) for each amino acid. This quantity is a key descriptor of the topological features of a protein’s specific conformation. Specifically, it is used to produce the protein representations through complex networks in a procedure is called proteotronicsAlfinito et al. (2008, 2015). Furthermore, it is commonly used to generate two-dimensional maps of the tertiary structure, such as contact maps Vendruscolo et al. (1997); Alfinito et al. (2010).
The coordination number allows to easily locate the amino acids inside the protein, i.e. the lower this number, the more superficial is the amino acid. On the other side, the more internal is the amino acid , the less effect the solvent has on it. Based on this approach, permittivity values are assigned so that highly connected regions correspond to lower permittivity, while less connected regions are assigned higher values. This strategy provides a more accurate estimate of buried versus exposed regions, thereby addressing limitations reported in previous studies Li et al. (2013).
@@
To define nearest neighbors, we model the protein as a complex network in which nodes correspond to the centroids of amino acids Alfinito et al. (2015, 2011, 2010). For each node, nearest neighbors are defined as those nodes located within a distance less than $R_{C}$, the interaction radius. The choice of $R_{C}$ determines the number of nearest neighbors: very small values (< 4 Å) produce a disconnected network, where most nodes have only 0–1 neighbors, whereas very large values result in a fully connected network, with each amino acid linked to all others Alfinito et al. (2008, 2015, 2011). Neither of these extremes is of interest. In this study, we set $R_{C}=6$ Å which captures the geometric regularity of proteins Simonson (2003); Alfinito et al. (2008, 2010) and yields an average of approximately five nearest neighbors per node.
@@
Finally, an appropriate distribution function must be selected. In principle, different choices may lead to significantly different outcomes. To address this issue, we evaluated several candidate functions, seeking those that yield higher permittivity values for elongated proteins than for spherical ones, while ensuring consistency across structurally similar proteins. These functions are formulated using nearest-neighbor descriptors, including the maximum value ($L_{\rm max}$) and the mean value ($\langle L\rangle$) computed over the entire protein. Overall, only minor differences were observed among the test-functions, particularly when normalized data were considered.
Finally, we compare our data with those obtained from a macroscopic procedure that involves calculating the permittivity of the protein’s electric dipole . This type of investigation is allowed by the Protein Dipole Moment Server Felder et al. (2007)developed in 2007 by Felder, Prilusky, and Sussman, which calculates the dipole moment of most protein entries in the protein data bank (PDB)Berman et al. (2000). Remarkably, the agreement is quite good and suggests further investigations .
In Fig1 a schematic view of the entire procedure is resumed.
Figure 1: Summary of the research described in this article. a. The 3D structure of the protein is obtained from the Protein Data Bank; b. As part of the proteotronic approach, the structure is mapped onto a complex network; c. The permittivity is calculated from the coordination number; d. The protein dipole, calculated as in Ref.Felder et al., 2007, is used to produce an alternative value of the protein permittivity; e. Finally, the data obtained from the two different approaches are compared. To make the figure we computed charges by the tool PDB2PQR (https://www.poissonboltzmann.org) with AMBER force field, and visualised the corresponding dipole momentum inside PyMol ( https://www.pymol.org ).
II Model and Results
In this section, we demonstrate two different procedures: one microscopic, inspired by continuum models, and the other macroscopic, based on a classical electrodynamics approach.
The microscopic approach is implemented using some test functions, as described in detail in subsection A, while the macroscopic approach, subsection B, demonstrates how to distinguish the dielectric response of a hydrated and dry protein. Finally, we compare the results obtained with the two methods.
II.1 Continuous model: Test-functions
@@
We consider a pool of test-functions employed to assign a local permittivity value, $\epsilon(n)$, to each $n$-th amino acid. [^1] Hereafter, the term permittivity refers to the relative permittivity, i.e., the permittivity normalized to the vacuum permittivity. This quantity depends on both the intrinsic (dry) permittivity, $\epsilon_{i}(n)$, and the amino acid’s position within the native protein structure. More specifically, $\epsilon(n)$ varies between a maximum value corresponding to that of the solvent under standard temperature and pressure (STP) conditions and $\epsilon_{i}(n)$, obtained using the polarizability values calculated in Song (2002). More details are given in Appendix B. We use three quite standard test-functions, namely, linear, sigmoidal and exponential, whose sharpness can be tuned via a control parameter (Appendix A). Increasing the sharpness enhances the contrast between regions characterized by high and low permittivity, facilitating the identification of hydrated (“wet”) regions associated with larger $\epsilon(n)$ values. The distinction between an internal low-permittivity region and an external high-permittivity region underlies many theoretical and computational studies Simonson and Brooks (1996); Simonson (2003, 2008); Warshel et al. (2006), as the effective dielectric response critically depends on this partitioning.
@@
[…]
All test-functions show a comparable trend, which suggests that it is not the chosen function itself that is important, but rather the permittivity distribution based on the coordination number. For illustration, Fig. 2 reports the dataset-averaged permittivity obtained using the linear distribution function, shown separately for elongated and spherical proteins. Further details on the model functions are provided in Appendix A.
@@
II.2 Macroscopic susceptibility
To recall, within the framework of classical electrodynamics, the electrical polarizability of a collection of weakly interacting dipoles is determined by the ensemble-averaged dipole orientation induced by an external electric field, $E$. Each dipole rotates to reduce the potential energy, say $U=-\bm{p_{0}}\cdot\bm{E}$. On the other hand, thermal fluctuations work against perfect alignment, such that for a freely rotating dipole, the higher the temperature, the less likely it is to remain antiparallel to the applied field $\bm{E}$.
The derivation follows from the thermal angular distribution function describing dipole orientations
@@ II.2 Macroscopic susceptibility
Proteins generally possess a not-negligible electric dipole momentSimonson and Brooks (1996); DE RIENZO et al. (2000); Simonson (2003).
Functioning proteins are usually hydrated protein, since the protein native structure itself arises from hydration through the folding process, which typically positions hydrophilic residues toward the solvent-exposed exterior and hydrophobic residues within the protein core. Water molecules near the surface dynamically reorient in response to local electrostatic fields, thereby reducing the internal electric interactions generated by charged and polar residues and contributing to structural stability. Consequently, the dipole moment reported in Ref. Felder et al., 2007 reflects this solvent-mediated charge organization, depends on single residue charges and dipoles and on the global protein shape. Finally, hydrated proteins cannot be regarded as freely rotating entities, as the surrounding solvent constrains their orientation in the presence of external electric fields. As a consequence, when considering an ensemble of hydrated proteins, the macroscopic response to an external electric field, permittivity, is smaller than that produced by the same ensemble when considered free of rotating.
In contrast, the dipole moment of the dry protein differs due to the absence of solvent-induced polarization and screening effects.
The dipole moment of a dry protein, denoted as $p_{0}$, can be estimated following the approach proposed by Song Song (2002). In that paper, the intrinsic polarizabilities of amino acids in proteins were evaluated using molecular dynamics simulations. The term intrinsic refers to a condition in which the dipole moment of each amino acid is assumed to be unaffected by external influences (e.g., solvent effects) and by interactions with other dipoles within the protein Song (2002). This approximation corresponds to the limiting case of a dry protein, where the dipole moments of individual amino acids are sufficiently small to be considered nearly independent Simonson (2003), and solvent-induced polarization is absent. Within this framework, the relative dielectric constant of a protein can be evaluated as $k_{0}=\langle\epsilon(n)\rangle$, where, for each amino acid, $\epsilon(n)=\epsilon_{i}$, i.e., the intrinsic dielectric constant obtained from Ref. Song, 2002 (see Appendix A). The resulting value of $k_{0}$ is more or less the same across the selected proteins, say, $k_{0}\approx 1.5$ , in fine agreement with the outcomes of Ref. Simonson, 2003; Seyedi and Matyushov, 2018 (see also Appendix B for details).
As a final comment, a collection of dry proteins may be regarded as freely orientable in an external electric field, due to the absence of solvent-induced friction. Consequently, the dielectric response of the system,i.e., the dry susceptibility, $\chi_{0}$, can be described by Eq. 4.
Otherwise, solvent both reduces the ability of the dipole of a hydrated protein to orient in external electric fields and modifies its value from $p_{0}$ to $p$. In other terms, $p=p(p_{0})$, which allows to calculate the expectation value of polarization under hydrated conditions as follows:
$\langle p^{\parallel}(T)\rangle=\int p\,\cos\theta\,\mathcal{P}(T)\,d\cos\theta.$ (5)
@@
Finally, we note that Eq. 8, yields an effective macroscopic susceptibility scaling as $\chi^{H}\propto p\,\Omega^{-q}\equiv\chi^{H}(q)$, with $q=1/2$. To assess whether the choice $q=1/2$ optimizes the agreement with the continuous microscopic models, Fig. 4 compares the previously introduced linear model with $\chi^{H}(q)$ evaluated for $q=0$, $0.5$, and $1$ , i.e. the original permittivity $\chi$. Within the adopted approximations, the closest agreement is observed for $q=1/2$.
@@
III Discussions and conclusions
The calculation and definition of permittivity in proteins remain subjects of ongoing debate and are known to depend sensitively on the adopted theoretical and computational framework (see, e.g., Ref. Warshel et al., 2006). On the other hand, this topic is of outmost importance for both fundamental and applicative researches, as for example, the development of biosensors or, in general, medical-oriented studiesTorculas et al. (2016); Valagiannopoulos and Sihvola (2021); Huang et al. (2024); Di Franco et al. (2025), therefore research is continuously evolving and exploring different strategies for interpreting and calculating the of main characteristics of the phenomenon Warshel et al. (2006); Li et al. (2013); Sauer et al. (2024); Seyedi and Matyushov (2018); Amin and Küpper (2020); Hartl et al. (2022); Sauer et al. (2024); Tenenbaum (2024); Agelii et al. (2025).
In this paper, we propose a computationally efficient approach for estimating the effective macroscopic susceptibility of hydrated proteins. This formulation is motivated by the need to describe the orientational response of macromolecules with restricted rotational mobility arising from their interaction with a polar solvent.
The investigation illustrated here is divided into two steps. In the first, we propose a microscopic study of the permittivity, $k$, of hydrated proteins, analyzing the interaction between individual amino acids and the solvent. This is done using a continuous distribution model of the permittivity of individual amino acids, assigning a higher value to those closest to the solvent molecules. This description simply translates the idea that while there is an intrinsic (dry) permittivity of the amino acids, the same observed in the inner part of the protein, whose effective value is provided by the simultaneous presence of the amino acids and the solvent. In electrical engineering, this is equivalent to a parallel circuit between the solvent and the biological material and is in perfect agreement with the description used in more refined procedures, such as, for example, Ref.Simonson and Brooks, 1996. The measured permittivity of a hydrated protein is derived from the effective dielectric constant. On the other hand, the permittivity of the same protein, in the absence of water, is simply derived from the intrinsic permittivity.
@@ III Discussions and conclusions
The second step involves the macroscopic analysis of permittivity, starting from the whole protein intrinsic dipole. In classical electrodynamics, in the presence of an intrinsic dipole, its orientation provides the main contribution to the shielding of electric fields: This response is quantified macroscopically in terms of the material’s susceptibility and, ultimately, its permittivity. In particular, a system of freely rotating dipoles yields the susceptibility given by Eq. 3. This expression is not directly applicable to hydrated proteins (dipole moment $p$), as it leads to susceptibility values significantly larger than those reported experimentally (see Table 1). The discrepancy arises from the fact that proteins in aqueous environments are not freely orientable: their rotational dynamics are constrained by solvent interactions (i.e., viscosity), which in turn limit their dielectric response. By contrast, the susceptibility evaluated using Eq. 8 produces values qualitatively consistent with the expected range Simonson (2003)(see Table 1). Furthermore, the susceptibility associated with dry proteins (dipole moment $p_{0}$), denoted as $\chi_{0}$, falls within the range typically attributed to the internal regions of proteins Simonson (2003). We emphasize that the difference between $p$ and $p_{0}$ does not primarily originate from structural modifications of the three-dimensional conformation, but rather from changes in polarizability. The dipole moment $p$ characterizes a hydrated protein, where interfacial water molecules dynamically reorient and influence the distribution of surface charges, thereby altering the dielectric response, particularly for solvent-exposed residues. In contrast, $p_{0}$ corresponds to the limiting case of a dry protein, in which the native structure is ideally preserved while solvent-induced polarization and screening effects are absent. Within this interpretation, $p$ represents a hydration-dependent state derived from $p_{0}$, where the dominant variation concerns the effective permittivity. Accordingly, $p$ may be regarded as a function of $p_{0}$, and its expectation value can be estimated using the orientational probability distribution associated with the freely rotating dipole $p_{0}$. The results obtained using the two independent procedures show qualitative consistency, supporting the validity of the proposed framework. The comparison of results from two different computational and design approaches is in itself a step forward in understanding the multiple aspects of permittivity in proteins, an issue which is well described by several authors and, in particular, in Refs.Schutz and Warshel, 2001; Warshel et al., 2006. Nevertheless, further refinements are possible, particularly regarding the selection of test functions, the choice of the interaction radius, $R_{C}$ and the evaluation of dipole moments. Finally, we note that in physiological and experimental conditions proteins are generally not freely mobile. This observation suggests that additional investigation is warranted to clarify the definition and interpretation of electrical susceptibility in constrained biological environments.
The proposed procedure is not intended to provide a complete or fully rigorous description. Rather, it offers a practical strategy that may complement more detailed molecular dynamics approaches, which, despite their accuracy, have not yet produced a comprehensive reference dataset of dielectric properties across diverse protein classes.